OpenStudy (anonymous):

What is the simplified form of (square root of 48n to the 9th)

OpenStudy (mertsj):

\[\sqrt{48n^9}=\sqrt{16n^8(3)(n)}=\sqrt{16}\sqrt{n^8}\sqrt{3n}=4n^4\sqrt{3n}\]